# Ctrl-Crash: Controllable Diffusion for Realistic Car Crashes

### v1

@@ Appendix G Training and Implementation Details
Model Architecture.
Ctrl-Crash builds on the Stable Video Diffusion (SVD) framework as the base image-to-video generation model. In the first stage, we fine-tune the SVD model on curated crash and driving clips from the MM-AU dataset (see Section 3.2). In the second stage, we freeze the base model and train a ControlNet module to integrate spatial (bounding box control frames) and semantic (crash type) inputs via additional encoder and cross-attention layers. All control signals are processed using the same pretrained VAE encoder used by SVD. The number of parameters for each model (and each sub-module) is given in Table 6.
Training Setup.
We use the AdamW optimizer with a learning rate of $4\times 10^{-5}$ and batch size of $1$. The first-stage fine-tuning of the base SVD model is performed for $101$k steps, using an MSE loss in latent space. The second-stage ControlNet training is run for $31$k steps, with conditioning dropout applied as described in Section 4.4. We use mixed precision during training by setting weights and inputs to fp16 for non-trainable (frozen) parts of the model (i.e., VAE encoder, VAE decoder, CLIP encoder) and keep the trainable parts at fp32 to reduce memory usage.Training is performed on 4 NVIDIA 80GB A100 GPUs over approximately 2 weeks for both stages combined. All models are implemented in PyTorch using the Hugging Face accelerate library as a base.
Inference and Guidance Parameters.
@@
Dataset Splits.
We curate a training set of $7,500$ clips from MM-AU after filtering. A held-out validation set of $900$ clips is used for evaluation. All reported metrics in the main paper (Tables 1, 2, 3) are computed on generated samples from the held-out validation set.
@@
Appendix H Datasets
Dataset annotations may be obtained by following instructions from our open-source code base (find link to code repo on our project website: https://anthonygosselin.github.io/Ctrl-Crash-ProjectPage/).
@@ Appendix H Datasets
Ctrl-Crash Types DADA2000 Crash Types
1 - ego-only 13, 14, 15, 16, 17, 18, 61, 62
2 - ego/vehicle 1-12
3 - vehicle-only 19-37, 39, 41, 42, 44
4 - vehicle/vehicle 38, 40, 43, 45-51
@@
For training and evaluation, we split all videos into 25-frame clips. The MM-AU dataset gives exact frame labeling to indicate when the accident occurs and when abnormal driving starts and ends before and after the accident. For each video we sample a 25-frame video containing the accident frame and label it according to the accident type. Then we sample, if possible, a 25-frame video containing only "normal" driving (i.e., with no overlap with abnormal or accident labeled frames) and label these clips with a crash type "no crash". Ultimately this yields $6,927$ clips containing a crash and $1,964$ clips without a crash (see Table 8 for number of clips for each class during training).
Table 8: Number of clips per crash type in training set
Crash Type Number of Training Samples
0 - no crash $1745$
1 - ego-only $267$
2 - ego/vehicle $3182$
3 - vehicle-only $577$
4 - vehicle/vehicle $2168$
@@ Appendix I User Survey
Please help us get some human feedback on a new video generation method. We’re asking you to rank 5 sets of videos. Should take less than 5 min, and you’ll see why AI won’t take over anything anytime soon. Instructions: For each of these 5 questions, we will show you 3 short video clips that have ALL been AI-generated. We will ask you: Does each video depict a crash? Rank the 3 videos by highest physical accuracy/plausibility Rank the 3 videos by highest visual fidelity (aka are they nice to look at)
Important: Please rank the videos relative to one another, i.e. "best" means best of the 3.
Each of the 5 accident types (with 3 videos each, labeled A,B,C, randomly shuffled from AVD2, DriveGen, ours) were presented like in Fig.14. The 5 accident types we selected were: head-on collision, t-bone, rear-ending, dangerous overtaking, and loosing control of the vehivle and going off the road. The ranking was implemented as shown in Fig.15 by forced-choice.
As described in the main body of the paper, we used the Friedman test (from the Python package scipy.stats.friedmanchisquare with $\alpha\leq 0.01$ and found $p=0.0001$. For the Nemenyi Post-Hoc analysis, we used a critical value of $critical\_value=3.64$ for an alpha value of $\alpha=0.01$. The critical difference (cut-off value for significance) was then calculated as $cd=critical\_value*\sqrt{k*(k+1)/(12*num\_participants)}$ with $num\_participants=40$. We obtained significant differences between the pairs: Ours - AVD2 and Ours - DriveGen, with our method consistenly ranking higher than either. We did non obtain significant differences in user preferences between the pair: AVD2 - DriveGen, which means that users did not have a significant preference between the two.
@@
Figure 15: User Survey Screenshot 2, showing the evaluation questions for each batch of 3 videos. Users were forced to rank all videos and to only use each rank once (i.e. it is not possible to submit the form when more than one video in each batch has the same rank).

### v2

@@ Appendix G Training and Implementation Details
Model Architecture
Ctrl-Crash builds on the Stable Video Diffusion (SVD) framework as the base image-to-video generation model. In the first stage, we fine-tune the SVD model on curated crash and driving clips from the MM-AU dataset. In the second stage, we freeze the base model and train a ControlNet module to integrate spatial (bounding box control frames) and semantic (crash type) inputs via additional encoder and cross-attention layers. All control signals are processed using the same pretrained VAE encoder used by SVD. The number of parameters for each model (and each sub-module) is given in Table 6.
Training Setup
We use the AdamW optimizer with a constant learning rate of $4\times 10^{-5}$ and batch size of $1$. Extensive hyperparameter tuning was intentionally avoided, as our focus is on validating the method’s effectiveness rather than maximizing performance through optimization. The first-stage fine-tuning of the base SVD model is performed for 101101k steps, using an MSE loss in latent space. The second-stage ControlNet training is run for 3131k steps, with conditioning dropout applied as described in Section 3.4 (main paper). We use mixed precision during training by setting weights and inputs to fp16 for non-trainable (frozen) parts of the model (i.e., VAE encoder, VAE decoder, CLIP encoder) and keep the trainable parts at fp32 to reduce memory usage. Training is performed on 4 NVIDIA 80GB A100 GPUs over approximately 2 weeks for both stages combined. All models are implemented in PyTorch using the Hugging Face accelerate library as a base.
Inference and Guidance Parameters
@@
Dataset Splits
We curate a training set of $7,500$ clips from MM-AU after filtering. A held-out validation set of $900$ clips is used for evaluation. All reported metrics in the main paper (Tables 1, 2) are computed on generated samples from the held-out validation set.
@@
Appendix H Dataset
Dataset annotations will be made publicly available and it will be possible to obtain them by following instructions from our open-source code base. We include a subset of the annotations in the supplementary ZIP file for reference, as the full set would exceed the size limit for submission.
@@ Appendix H Dataset
Crash Type DADA2000 Crash Types # of Training Samples
0 - no crash N/A $1745$
1 - ego-only 13, 14, 15, 16, 17, 18, 61, 62 $267$
2 - ego/vehicle 1-12 $3182$
3 - vehicle-only 19-37, 39, 41, 42, 44 $577$
4 - vehicle/vehicle 38, 40, 43, 45-51 $2168$
@@
For training and evaluation, we split all videos into 25-frame clips. The MM-AU dataset gives exact frame labeling to indicate when the accident occurs and when abnormal driving starts and ends before and after the accident. For each video we sample a 25-frame video containing the accident frame and label it according to the accident type. Then we sample, if possible, a 25-frame video containing only ”normal” driving (i.e., with no overlap with abnormal or accident labeled frames) and label these clips with a crash type ”no crash”. Ultimately this yields $6,927$ clips containing a crash and $1,964$ clips without a crash (see Table 7 for number of clips for each class during training).
@@ Appendix I User Survey
(a) Visual preference comparison in generated videos.
(b) Realism preference comparison in physical appearance.
Figure 15: User study results comparing generated videos from AVD2, DrivingGen and Ctrl-Crash with 40 participants across 5 crash types: Participants exhibit a strong preference for Ctrl-Crash generated videos, citing superior visual quality and physical realism.
We conducted a brief user study with $n=21$ participants who were asked to rank $k=3$ videos (from Ctrl-Crash, AVD2, and DrivingGen) across 5 different crash scenarios each. The participants consist of students from our lab who are not associated with the project in any way. No further demographic data was collected. The users were asked to rank the 3 videos in each of the 5 situations by (a) physical plausibility and (b) visual fidelity from best to worst. The users were required to choose a best/medium/worst video in each question and visual/physical category. We used the non-parametric Friedman test [38] to determine with $p\leq 0.01$ that there is a method that is consistently ranked higher than the others (Ctrl-Crash). We further used the Nemenyi post-hoc analysis [39] to find that with $p\leq 0.01$, our method consistently outperforms both AVD2 and DrivingGen in both physical realism and visual fidelity (see Figure 15).
@@
Please help us get some human feedback on a new video generation method. We’re asking you to rank 5 sets of videos. Should take less than 5 min, and you’ll see why AI won’t take over anything anytime soon. Instructions: For each of these 5 questions, we will show you 3 short video clips that have ALL been AI-generated. We will ask you: 1. Does each video depict a crash? 2. Rank the 3 videos by highest physical accuracy/plausibility 3. Rank the 3 videos by highest visual fidelity (aka are they nice to look at)
Important: Please rank the videos relative to one another, i.e. ”best” means best of the 3.
Each of the 5 accident types (with 3 videos each, labeled A,B,C, randomly shuffled from AVD2, DriveGen, ours) were presented like in Figure 17. The 5 accident types we selected were: head-on collision, t-bone, rear-ending, dangerous overtaking, and loosing control of the vehivle and going off the road. The ranking was implemented as shown in Figure 17 by forced-choice.
@@
Figure 17: User Survey Screenshot 2, showing the evaluation questions for each batch of 3 videos. Users were forced to rank all videos and to only use each rank once (i.e., it is not possible to submit the form when more than one video in each batch has the same rank).